# Manipulation Priors Learned From Egocentric Videos

### v1

@@ Abstract
Large-scale egocentric video datasets capture diverse human activities across a wide range of scenarios, offering rich and detailed insights into how humans interact with objects, especially those that require fine-grained dexterous control. Such complex, dexterous skills with precise controls are also crucial for many robotic manipulation tasks, yet are often insufficiently addressed by traditional data-driven approaches to robotic manipulation. To address this gap, we leverage manipulation priors learned from large-scale egocentric videos to improve policy learning for dexterous robotic manipulation tasks. We present MAPLE, a novel method for dexterous robotic manipulation that exploits rich manipulation priors to enable efficient policy learning and better performance on diverse, complex manipulation tasks. Specifically, we predict hand-object contact points and detailed hand poses at the moment of hand-object contact and use the learned features to train policies for downstream manipulation tasks. Experimental results demonstrate the effectiveness of MAPLE across existing simulation benchmarks, as well as a newly designed set of challenging simulation tasks, which require fine-grained object control and complex dexterous skills. Our newly designed simulation environments address the shortage of dexterous manipulation benchmarks in the literature. The benefits of MAPLE are further highlighted in real-world experiments using a dexterous robotic hand, while simultaneous evaluation across both simulation and real-world experiments has often been underexplored in prior work. Our code, trained model, and the new manipulation benchmark suite will be made publicly available on https://algvr.com/maple/.
Figure 1: MAPLE: Encoding Dexterous Robotic Manipulation Priors Learned From Egocentric Videos. We present MAPLE, a framework that learns dexterous manipulation priors from egocentric videos and produces features well-suited for downstream dexterous robotic manipulation tasks. Experiments in both simulation and real-world settings demonstrate that MAPLE enables efficient policy learning and improves generalization across various tasks.
footnotetext:  Correspondence to agavryushin@ethz.ch
@@ 1 Introduction
As robotic systems become increasingly deployed across various domains, dexterous manipulation emerges as a critical capability, going beyond basic automation to enable robots to augment human abilities and function autonomously in dynamic, real-world environments. Consider a household robot tasked with preparing a simple meal, such as spaghetti with tomato sauce. To complete the task, the robot must interact with various kitchen objects–locating a pan, filling it with water, and placing it on the stove to boil. In such scenarios, the robot encounters a wide range of everyday objects, many of which require complex dexterous manipulation skills. Robust dexterous manipulation priors are essential for successfully performing these tasks in the real world. What is then the best source of information for teaching robots dexterous manipulation skills?
@@
Recently, the computer vision community has shown a growing interest in the egocentric perspective [15, 9, 64, 16, 26], as it gives unique insights into human interactions with objects. Large-scale egocentric datasets, in particular, capture the complex, dexterous dynamics of hand-object interactions across a wide range of object categories. It thus suggests itself that these datasets are well-suited for learning object manipulation priors directly from natural human-object interactions. Several recent studies [40, 32, 53, 55] have explored this direction by training visual representations on large collections of human egocentric videos. However, these approaches generally lack a specific focus on dexterous manipulation and often struggle to generalize, particularly in human-centric environments where fine-grained control is essential [46, 3, 11, 47].
@@
Why has learning generalizable dexterous manipulation priors remained so challenging? While general-purpose representations trained on large datasets capture rich scene-level details [45, 4], they lack a specific focus on manipulation-relevant cues, making them suboptimal. Other self-supervised learning approaches (e.g. VIP [31], trained with contrastive losses) struggle to effectively extract dexterous manipulation information from egocentric videos, especially due to the complex and cluttered nature of real-world scenarios depicted in these videos. This raises a key question: what specific information should a visual encoder designed for object manipulation learn to extract? We hypothesize that low-level interaction cues, specifically object contact points and grasping hand poses, could provide a strong prior for downstream dexterous manipulation tasks. To this end, we introduce MAPLE, a novel method that learns dexterous manipulation priors from egocentric videos for robotic manipulation.
Our key insight is to design learning objectives that guide representations to effectively capture dexterous manipulation knowledge. More specifically, we focus on training a visual encoder aimed at extracting where and how to interact with objects during manipulation. We use existing state-of-the-art methods to automatically extract supervision signals for training, greatly reducing the cost of dataset acquisition. Given a single image of the target object, MAPLE learns to predict hand-object contact points and the 3D hand pose at the moment of contact. The learned visual representations are then used as input to policy networks trained to control dexterous robot hands.
@@
We demonstrate that MAPLE enables the extraction of features from egocentric videos that are well-suited for robotic manipulation tasks requiring complex dexterous skills. We conduct experiments on established simulation-based benchmarks, specifically designed to evaluate dexterous manipulation skills. Additionally, we evaluate MAPLE on a newly designed set of challenging simulated manipulation tasks, which require more fine-grained object control and advanced dexterous skills. A comprehensive evaluation is performed across existing and newly proposed dexterous manipulation tasks, comparing MAPLE against general-purpose visual encoders and state-of-the-art methods. The benefits of MAPLE are further highlighted in real-world experiments using a dexterous robotic hand. Note that such an evaluation setting using both simulated and real-world experiments has been largely underexplored in prior work. We will publicly release our codebase, data extraction pipeline, extracted data, and the newly designed benchmark to support future research. In summary, our contributions are three-fold:
@@
  • We propose MAPLE, a novel visual encoder designed for dexterous manipulation tasks by learning to predict low-level cues for object manipulation.
@@
  • We propose a new set of custom-designed dexterous manipulation tasks in simulation that require more fine-grained object control.
@@
  • We demonstrate that MAPLE improves performance on downstream dexterous robotic manipulation tasks in both simulated environments and real-world settings, highlighting its improved generalization ability.
@@ 2 Related Work
Dexterous Manipulation. Dexterous manipulation describes the manipulation of objects with multi-fingered hands. Due to the high-dimensional action space and complex physical interactions, it is one of the most challenging tasks in robotics. Traditional approaches in dexterous manipulation typically rely on analytical grasp planning methods that use models [6] or optimization [61, 63]. These grasp proposals are then executed in an open loop, feed-forward fashion using a suitable task-space controller. While this modular approach allows for more flexibility, it faces challenges when dealing with partially observed environments or when adaptability and tight coupling of vision and control are necessary. Therefore, more recent work has investigated the use of data-driven end-to-end techniques such as reinforcement learning to learn manipulation policies [69, 68, 43, 60]. Reinforcement learning provides a powerful tool to learn robust dexterous manipulation policies, but its performance heavily relies on simulation accuracy as well as proper reward formulations, which might be challenging to formulate. This has sparked research interest that led to the advent of imitation learning [5, 70, 22, 28], which directly leverages human demonstrations as a supervision signal. While most of these methods are able to learn meaningful manipulation policies, their generalization across different task is still limited – often relying on a large amount of task-specific trajectories in order to be used for new tasks.
@@
Learning From Egocentric Videos. Large-scale human-recorded visual observations have been widely used to extract embeddings or motion priors for robotic manipulation. Existing methods can largely be divided into implicit and explicit methods.
@@
Implicit methods [40, 31, 35, 32] utilize large-scale contrastive learning to derive meaningful features from egocentric interaction data. Some works, such as [31, 32, 66] leverage large-scale video datasets to learn goal-conditioned value functions [31] or train masked autoencoders (MAE) [66, 32]. Additionally, methods like [40] integrate language embeddings by applying a contrastive loss across image and language modalities. However, due to their implicit nature, these methods can’t be directly used for robotic manipulation. Instead, their pre-trained (frozen) embeddings serve as inputs for reinforcement learning [66] or imitation learning [40] algorithms.
@@
Conversely, explicit methods [39, 1, 2, 21, 34] aim to directly learn interaction affordances from videos by using human interaction as an explicit supervision signal. These approaches typically predict interaction trajectories [39], contact heatmaps [34, 1, 21], or future contact points and wrist position [55]. While their output could be directly used within a task-specific controller, most methods use the final predictions or intermediate embeddings in policy learning due to the complexity of the manipulation task. Although explicit supervision helps learn a better-suited latent space, most existing approaches focus on high-level interactions and are primarily designed for simple two-finger manipulation. Furthermore, they often depend on human-annotated data, limiting scalability. In contrast, our method does not rely on any human-annotated data and adopts a more fine-grained formulation capturing detailed manipulation dynamics, well-suitable for dexterous tasks.
@@
Evaluation in Robotic Simulators. We categorized simulation benchmarks by the type of end-effector used. Benchmarks like MetaWorld [67], Franka Kitchen [14, 17], and Habitat 2.0 [56] focus on parallel jaw grippers, defining tasks such as object manipulation, cooking-related activities, and home environment rearrangement tasks. The DeepMind Control Suite [57] also includes basic object manipulation tasks in abstract environments. For three-finger grippers, TriFinger [65] provides a benchmark for object manipulation, particularly evaluating toy cube tasks in both real and simulated settings. Finally, dexterous hand benchmarks offer more complex manipulation challenges. DAPG [47] introduces four environments for a simulated Adroit [24] hand, while DexMV [44] includes tasks like pouring, object relocation, and target-based movement, requiring more fine-grained control. A key limitation across many benchmarks is their reliance on low-DoF grippers, restricting their ability to evaluate policies for fine-grained manipulation. This motivates us to design a more comprehensive dexterous manipulation benchmark.
@@
Evaluation in Real World. In the robotics community, imitation learning is becoming a standard approach to performing dexterous manipulation tasks [5, 70, 12, 7], even driving the development of foundation model-style generalist robot policies that can be fine-tuned with target data from small samples [58, 23]. This is because it enables learning from task demonstrations without the need for simulation environments or task-specific rewards. However, these models have been primarily trained and deployed on low-DoF grippers. Furthermore, for dexterous manipulators, there are no standard datasets due to the high variation between designs. This motivates us to design dexterous manipulation tasks in the real world, reproducible with a commercially available dexterous hand [36].
@@ 3 MAPLE
We present Encoding Manipulation Priors Learned from Egocentric Videos (MAPLE), an approach for learning manipulation priors from egocentric videos to enable dexterous robotic manipulation. The MAPLE pipeline involves solving two key problems: how to model manipulation priors and extract corresponding labels from egocentric videos, and how to learn these priors from the extracted labels. Figure 2 illustrates the entire pipeline.
Figure 3: Extracting Contact Points and Hand Poses From Egocentric Videos. We first extract contact points and hand poses from the contact frame, and use a point tracker to backtrack the contact locations on the prediction frame.
3.1 Extracting Manipulation Labels From Videos
MAPLE aims to learn features useful for dexterous manipulation from large, publicly available egocentric video datasets. To effectively represent manipulation priors, we consider the contact points between hands and objects, along with the corresponding 3D hand poses at the moment of contact. We use off-the-shelf tools to automatically parse these hand-object interactions and extract the relevant training labels.
Specifically, we define a contact frame $f_{c}$, which marks the moment when the hand first makes contact with the object. We also define a prediction frame $f_{p}$, occurring earlier in the video than $f_{c}$, when the hand is sufficiently distant from the object it is about to interact with. From the contact frame $f_{c}$, we extract the hand pose vector $\mathcal{H}_{c}$ representing the interacting hand. Using $\mathcal{H}_{c}$, we further extract the contact points $pt_{c}$ and then track back to $f_{p}$ to obtain the future contact points $pt_{p}$ on the prediction frame. Figure 3 provides a visualization of our label extraction procedure.
@@ 3.1 Extracting Manipulation Labels From Videos
Extract Contact Supervision From the Contact Frame. To locate moments of contact, we process all frames in the dataset using an off-the-shelf hand-object contact segmentation model  [10] and consider the first contact frame as $f_{c}$. To avoid extracting identical $f_{p}$ from neighboring contact frames, we only process contact frames with no other $f_{c}$ preceding in the last $\delta_{min}=90$ frames.
We define contact points $pt_{c}$ as the points on the object where the fingertips of the thumb and index finger first make contact with the object. For each contact frame, we use a pre-trained hand pose estimation model [42] to extract the pose vector of the interacting hand, resulting in MANO [50] hand poses $\mathcal{H}_{c}\in\mathbb{R}^{21\times 3}$. The extracted $\mathcal{H}_{c}$ are forwarded through a MANO layer to obtain 2D fingertip keypoints. We then project the thumb and index fingertip keypoints onto the object mask to obtain $pt_{c}$. When calculating the position of the projected fingertip points, we use a slightly smaller object mask from which a portion of the boundary has been erased by $n_{e}=12$ binary erosion [52] operations. This helps when finding the contact points in the prediction frame in the next step. Details including visualizations and failure analysis are provided in Supp. Mat.
@@
Extract Supervision From the Prediction Frame. To identify the prediction frame $f_{p}$, we run a point tracker [48] to project the contact points backward in video time until we find the first frame where the hand mask does not yet intersect the object mask. Using smaller object masks for contact point identification helps prevent the point tracker from tracking the fingers of the manipulating hand instead of the object. When checking for the intersection, we expand the object mask by $n_{d}=75$ binary dilation [52] iterations to ensure a sufficient distance from the object to the hand. This helps make the contact point regression less trivial than simply locating the object immediately next to the hand. Prior work [55, 1] randomly samples contact points from the intersection of the bounding boxes induced by the hand and the manipulated object, which is highly prone to erroneous results, such as when points are sampled from the background as the hand is touching a concave object. In contrast, we project the fingertips to the manipulated objects’ masks to extract precise contact information.
@@
Tokenizing Hand Poses. Learning to predict hand poses associated with grasping a target object is a challenging task, due to the requirement of conforming to anatomical constraints and reasoning about a complex kinematic tree with a high number of joints. Furthermore, a large subset of the joint configuration space results in hand poses exhibiting self-penetration. We simplify the prediction task by reframing it as a classification task and letting the model select the hand pose among a limited list of “reasonably frequent” (i.e., observed in videos) and anatomically correct choices rather than regress the correct hand pose by itself. For that purpose, we train a hand pose tokenizer on hand poses extracted by HaMeR [42] from a randomly selected subset of the Ego4D dataset. We use the same Memcodes [33]-based tokenization scheme as the human pose tokenizer proposed in 4M [37], employing 8 codebooks of size 1024 each to represent a single MANO [50] hand pose. Qualitative reconstruction examples of the hand tokenizer can be found in the Supp. Mat.

### v2

@@ Abstract
Large-scale egocentric video datasets capture diverse human activities across a wide range of scenarios, offering rich and detailed insights into how humans interact with objects, especially those that require fine-grained dexterous control. Such complex, dexterous skills with precise controls are also crucial for many robotic manipulation tasks, which are often insufficiently addressed by traditional data-driven approaches to robotic manipulation. To address this gap, we leverage manipulation priors learned from large-scale egocentric videos to improve policy learning for dexterous robotic manipulation tasks. We present MAPLE, a novel method for dexterous robotic manipulation that exploits rich manipulation priors to enable efficient policy learning and better performance on diverse, complex manipulation tasks. Specifically, we predict hand-object contact points and detailed hand poses at the moment of contact and use the learned features to train policies for downstream manipulation tasks. Experimental results demonstrate the effectiveness of MAPLE across existing simulation benchmarks, as well as a newly designed set of challenging simulation tasks, which require fine-grained object control and complex dexterous skills. Our newly designed simulation environments address the shortage of dexterous manipulation benchmarks in the literature. The benefits of MAPLE are further highlighted in real-world experiments using a dexterous robotic hand, while simultaneous evaluation across both simulation and real-world experiments has often been underexplored in prior work. Our code, trained model, and the new manipulation benchmark suite will be made publicly available on https://algvr.com/maple/.
Figure 1: Encoding Dexterous Robotic Manipulation Priors Learned From Egocentric Videos. We present MAPLE, a framework that learns dexterous manipulation priors from egocentric videos and produces features well-suited for downstream dexterous robotic manipulation tasks. Experiments in both simulation and real-world settings demonstrate that MAPLE enables efficient policy learning and improves generalization across various tasks.
11footnotetext:  Correspondence to agavryushin@ethz.ch
@@ 1 Introduction
With robotic systems expanding into diverse applications, dexterous manipulation emerges as a critical capability to interact autonomously in dynamic, real-world environments. Considering a household robot tasked with preparing a simple meal, the robot must interact with various kitchen objects, such as selecting a pan and placing it on the stove. In such scenarios, the robot encounters a wide variety of everyday objects, many of which require complex dexterous manipulation skills. What is, then, the best source of information for teaching robots these skills?
@@
Large-scale egocentric datasets [Ego4D, EK100, wang2023holoassist, grauman2024ego, li2024egogen] can offer a compelling answer, as they provide a scalable data source and capture rich, natural examples of how humans interact with objects. This makes them well-suited for learning visuo-motor manipulation priors directly from real-world natural human-object interactions. Several recent studies [R3M, VC1, singh2024hand, HRP] have explored this direction by pretraining visual representations on large-scale human egocentric videos. However, these approaches generally lack the focus on dexterous manipulation and often struggle to generalize, particularly in human-centric environments where fine-grained control is essential [radosavovic2023robot, burns2023makes, dasari2023unbiased, DAPG].
@@
Why, then, is learning generalizable dexterous manipulation priors still so challenging? While general-purpose representations pre-trained on large datasets capture rich scene-level details [CLIP, DINO, DINOv3], they lack a specific focus on manipulation-relevant cues, making them suboptimal. Other self-supervised learning approaches (e.g. R3M [R3M], trained with contrastive losses) struggle to effectively extract dexterous manipulation information from egocentric videos, especially due to the complex and cluttered nature of real-world scenarios depicted in these videos. This raises a key question: what specific information should a visual encoder designed for object manipulation learn to extract? We argue that this encoder should be trained to reason about where to interact with an object and how to do so. More concretely, we hypothesize that object contact points and grasping hand poses provide strong priors for downstream dexterous manipulation tasks. To this end, we introduce MAPLE, a novel method that learns dexterous manipulation priors from egocentric videos for robotic manipulation. We develop a pipeline to automatically extract supervision signals from egocentric videos, enabling large-scale training without manual annotation. Given a single image of the target object, MAPLE learns to predict hand-object contact points and the 3D hand pose at the moment of contact. The learned visual representations are then used as input to visuo-motor policies that are trained to control dexterous robotic hands.
@@
Unlike previous work that predicts coarse affordance [HRP, kannan2023deft], we demonstrate that MAPLE learns transferable features by predicting fine-grained hand-object interaction cues, enabling diverse robotic embodiments to perform a wide range of manipulation tasks that require complex dexterous skills. We conduct experiments on established simulated benchmarks specifically designed to evaluate dexterous manipulation skills. Additionally, we introduce a new suite of challenging manipulation tasks in simulation that require fine-grained object control and advanced dexterous coordination. We first show that the features learned with MAPLE transfer beyond policy learning by evaluating them on a contact point prediction task. Then, we use them as input for learning visuomotor policies, performing a comprehensive simulation evaluation across four existing and four newly proposed dexterous manipulation tasks, comparing MAPLE against general-purpose visual encoders and state-of-the-art methods. The benefits of MAPLE are further highlighted in real-world experiments using a dexterous robotic hand. Such an evaluation setting using both simulation and real-world experiments has been largely underexplored in prior work. We will publicly release our codebase and data to support future research.
In summary, our contributions are:
@@
  • We propose MAPLE, a visual encoder for dexterous manipulation tasks, learned through a novel pretraining strategy that predicts interaction cues for object manipulation.
@@
  • We propose a new set of custom-designed dexterous manipulation tasks with new assets in simulation that require more fine-grained object control.
@@
  • We demonstrate that MAPLE significantly improves performance on downstream dexterous robotic manipulation tasks in both simulation environments and real-world settings as well as an egocentric contact point prediction task, highlighting its improved generalization ability.
@@ 2 Related Work
Dexterous Manipulation. Dexterous manipulation describes the manipulation of objects with multi-fingered hands. Due to the high dimensional action space and complex physical interactions, it is one of the most challenging tasks in robotics. Traditional approaches in dexterous manipulation typically rely on analytical grasp planning methods that use models [DGrasp] or optimization [turpin2023fast, wang2023dexgraspnet]. These grasp proposals are then executed in an open loop, feed-forward fashion using a suitable task-space controller. While this modular approach allows for more flexibility, it faces challenges when dealing with partially observed environments or when adaptability and the tight coupling of vision and control are necessary. Therefore, more recent work has investigated the use of data-driven end-to-end techniques such as reinforcement learning to learn manipulation policies [zhang2024artigrasp, zhang2024graspxl, qi2023hand]. Reinforcement learning provides a powerful tool to learn robust dexterous manipulation policies, but its performance heavily relies on simulation accuracy as well as proper reward formulations, which might be challenging to formulate. This has sparked research interest that has led to the advent of imitation learning [chi2023diffusion, zhao2023learning, kareer2024egomimic, lin2024learning, nava2025mimiconescalablemodelrecipe, bauer2025latentactiondiffusioncrossembodiment] which directly leverages human demonstrations as a supervision signal. While most of these methods are able to learn meaningful manipulation policies, their generalization across different tasks is still limited, and they often rely on a large amount of task-specific trajectories in order to be used for new tasks.
@@
Learning From Egocentric Videos. Large-scale human videos have been widely used to extract embeddings or motion priors for robotic manipulation. Existing methods can largely by divided into implicit and explicit methods.
@@
Implicit methods [R3M, Mandikal22DexVIP, VC1] use large-scale contrastive learning to derive meaningful features from egocentric interaction data. Some works, such as [VC1, xiao2022masked], leverage large-scale video datasets to train masked autoencoders (MAE). Methods like [R3M] integrate language embeddings by applying a contrastive loss across image and language modalities. However, due to their implicit nature, these methods cannot be directly used for robotic manipulation. Instead, their pre-trained (frozen) embeddings serve as inputs for reinforcement learning [xiao2022masked] or imitation learning [R3M] algorithms.
@@
Conversely, explicit methods [nagarajan2019grounded, VRB, Bharadhwaj23ZeroShotRobotManip, zeng2024mpi, Mandikal21GRAFF] aim to directly learn interaction affordances from videos by using human interaction as an explicit supervision signal. These approaches typically predict interaction trajectories [nagarajan2019grounded], contact heatmaps [Mandikal21GRAFF, VRB, zeng2024mpi], or future contact points and wrist positions [HRP]. While their output could be directly used within a task-specific controller, most methods use the final predictions or intermediate embeddings in policy learning due to the complexity of the manipulation task. Although explicit supervision helps learn a better-suited latent space, most existing approaches focus on high-level interactions and are primarily designed for simple two-finger manipulation. Furthermore, they often depend on human-annotated data, limiting scalability. In contrast, our method does not rely on any human-annotated data and adopts a more fine-grained formulation to capture detailed manipulation dynamics, well-suitable for dexterous manipulation.
@@
Evaluation in Robotic Simulators. We categorized simulation benchmarks by the type of end-effector used. Benchmarks like MetaWorld [MetaWorld], Franka Kitchen [D4RL, gupta2019relay], and Habitat 2.0 [Habitat2] focus on parallel jaw grippers, defining tasks such as object manipulation, cooking-related activities, and home environment rearrangement tasks. The DeepMind Control Suite [DMCS] includes basic object manipulation tasks in abstract environments. For three-finger grippers, TriFinger [TriFinger] provides a benchmark for object manipulation, particularly evaluating toy cube tasks in both real and simulated settings. Finally, dexterous hand benchmarks offer more complex manipulation challenges. DAPG [DAPG] introduces four environments for a simulated Adroit [AdroitHand] hand, while DexMV [DexMV] includes tasks like pouring, object relocation, and target-based movement, requiring more fine-grained control. A key limitation across these benchmarks is their overwhelming reliance on low-DoF grippers, leaving researchers few options to evaluate policies for fine-grained manipulation. This motivates us to design a more comprehensive dexterous manipulation benchmark.
@@
Evaluation in Real World. In the robotics community, imitation learning is becoming a standard approach to performing dexterous manipulation tasks [chi2023diffusion, zhao2023learning, dasari2024DiTpolicy], even driving the development of foundation model-style generalist robot policies that can be fine-tuned with target data from small samples [octomodelteam2024Octo, kim2024OpenVLA]. This is because doing so enables learning from task demonstrations without the need for simulation environments or task-specific rewards. However, these models have been primarily trained and deployed on low-DoF grippers. Furthermore, for dexterous manipulators, there are no standard datasets due to the high variation between designs. This served as motivation for our benchmark for dexterous manipulation tasks in the real world.
@@ 3 MAPLE
We aim to learn object manipulation skills from egocentric videos that capture natural human interactions with diverse everyday objects. To this end, we introduce MAPLE, a pipeline for learning manipulation priors. The pipeline addresses two key challenges: modeling manipulation priors (Sec. 3.1) by extracting supervision from videos (Sec. 3.2) and learning these priors for downstream dexterous control (Sec. 3.3). Sec. 3.4 details the training strategy. See Figure 2 for our method overview.
Our fully automated extraction pipeline produces approximately 82,100 training samples from the large-scale Ego4D [Ego4D] dataset, which we then use to train our visual encoder model for dexterous tasks.
Figure 3: Manipulation Prior Modeling. We learn manipulation priors by predicting future contact points and hand poses from an input frame. Training data are extracted by identifying a contact frame $f_{c}$ and a preceding prediction frame $f_{p}$, which is used as input. Contact points and hand poses are extracted from the contact frame, and a point tracker is used to back-project the contact locations onto the prediction frame.
3.1 Manipulation Prior Formulation
To effectively model manipulation priors, we focus on the moment of contact, which captures the transition from perception to physical interaction. At contact, the hand pose and contact points between hand and objects are jointly constrained, exposing the action’s affordance. This physically grounded state provides stable cues that are directly transferable to downstream manipulation learning.
Therefore, we consider both contact points and hand poses as the manipulation prior and formulate the following learning task. Given an input frame, our goal is to predict future contact points and hand poses upon contact. Training data are extracted by identifying a contact frame $f_{c}$ and a prediction frame $f_{p}$, which is used as input. We define $f_{c}$ as the moment when the hand first contacts the object and $f_{p}$ as an earlier frame, where the hand is sufficiently distant from the object. From $f_{c}$, we extract the interacting hand pose vector $\mathcal{H}_{c}$ and corresponding contact points $pt_{c}$, which are then back-projected to $f_{p}$ via point tracking to obtain future contact points $pt_{p}$ in the prediction frame. Figure 3 provides a visualization of this formulation.
3.2 Manipulation Supervision Extraction
@@
Identifying the Contact Frames. To locate moments of contact, we process all frames using the off-the-shelf hand-object contact segmentation model VISOR-HOS [EKVisor] and consider the first contact frame in a series as $f_{c}$.
@@
Extracting Contact Labels from the Contact Frame. We define the contact points $pt_{c}$ as the object surface points where the fingertips of the thumb and index finger first touch the object. For each contact frame, we use the pre-trained hand pose estimation model HaMeR [HaMeR] to extract the MANO [MANO] pose vector $\mathcal{H}_{c}\in\mathbb{R}^{21\times 3}$ of the interacting hand. $\mathcal{H}_{c}$ is forwarded through a MANO layer to obtain 2D fingertip keypoints. We then project the thumb and index fingertip keypoints onto the object mask to obtain $pt_{c}$. When calculating the projected fingertip positions, we use a slightly smaller object mask from which a portion of the boundary has been erased by binary erosion [BinaryDilation]. This also helps when finding the contact points in the prediction frame in the next step. Prior work [HRP, VRB] randomly samples contact points from the intersection of the hand bounding boxes and the manipulated object, which is highly error prone, such as when points are sampled from the background as the hand is touching an object on its visual boundary. In contrast, we project fingertip locations to the manipulated objects’ segmentation masks to extract precise contact information. Extracting the Prediction Frame and its Contact Labels. To identify the prediction frame $f_{p}$, we run a point tracker [SamPT] to project the contact points backward in time until reaching the first frame where the hand mask no longer intersects the object mask. Using smaller object masks for contact point identification helps prevent the point tracker from tracking the fingers of the manipulating hand instead of the object. When checking for the intersection, we expand the object mask using binary dilation [BinaryDilation] to ensure a sufficient distance from the object to the hand. This helps make the contact point regression less trivial than simply selecting the object next to the hand.
@@
Tokenizing Hand Poses. Learning to predict hand poses associated with grasping a target object is a challenging task, due to the need to conform to anatomical constraints and reason about a complex kinematic tree with a high number of joints. Furthermore, a large subset of the joint configuration space results in hand poses that exhibit self-penetration. We simplify the prediction task by reframing it as a classification task and letting the model select the hand pose among a limited list of “reasonably frequent” (i.e., observed in videos) and anatomically correct choices rather than regressing the correct hand pose by itself. For that purpose, we train a tokenizer on hand poses extracted by HaMeR [HaMeR] from a subset of the Ego4D dataset. We use the same Memcodes [mama2021nwt]-based tokenization scheme as the human pose tokenizer proposed in 4M [4m], employing 8 codebooks of size 1024 each to represent a single MANO [MANO] hand pose.
In Supp. Mat., we provide visualizations and an analysis of failure cases in the extracted manipulation supervision, the most common ones being false contact positives and samples with prediction frames still exhibiting contact. These residual errors reflect the large number of false positives generated by [EKVisor], which are largely eliminated by filtering in our pipeline. Qualitative reconstruction examples of the hand tokenizer can be found in the Supp. Mat.